Be sure to answer all parts.

Calculate the pOH of the following solutions at 25°C.

(a) [OH] = 5.00 × 10−11

pOH = (b) [OH]= 2.51 ×10−3

pOH =
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

formula used pon=log [on] 0 [on]= 5.00x10 pona - lug [5.00x10] por = 110-30 ] 3 [on] = 2.51x10-3 pox = -log [2.51x10-3] . p